OpenStudy (anonymous):

What is the slope of the line passing through the points (4, -2) and (-5, 4)?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (aravindg):

slope=y difference/x difference

OpenStudy (skullpatrol):

$$\Huge slope=\frac{\delta{x}}{\delta{y}}=\frac{y_2-y_1}{x_2-x_1}$$

OpenStudy (anonymous):

slope can be found using the formula m= (Y2-Y1)/(X2-X1) m is the slope of the line the order of which is 1 and 2 does not matter :P for (4,-2) and (-5,4) that would be m=(4-(-2))/(-5)-4) = (4+2)/(-5-4) = 6/-9 = -2/3 the slope of the line is {-2/3} hope it helped good luck :P